# Enhance Lifelong Model Editing with Continuous Data-Adapter Association

### v1

@@ Abstract
Large language models (LLMs) require model editing to efficiently update specific knowledge within them and avoid factual errors. Most model editing methods are solely designed for single-time use and lead to a significant forgetting effect after sequential edits over time, referred to as lifelong editing. Current approaches manage sequential edits by freezing original parameters and allocating new adapters for each knowledge modification. However, these methods lack robustness to minor input variations. To address this challenge, we propose ELDER, Enhancing Lifelong moDel Editing with mixtuRe of Low-Rank Adapter (LoRA). ELDER is an adaptive approach that integrates multiple LoRAs through a router network. It learns to create a continuous and smooth association between data and adapters, thereby enhancing robustness and generalization to semantically equivalent inputs. Additionally, we introduce a novel loss to help learn associations between adapter allocations and edit semantics. A deferral mechanism is also proposed to retain the original LLM capabilities post-edit. Extensive experiments on GPT-2 XL and LLaMA2-7B demonstrate that ELDER effectively edits models in the lifelong setting and exhibits strong scalability, while retaining LLM’s general abilities on downstream tasks.
Enhance Lifelong Model Editing with Continuous Data-Adapter Association
Jiaang Li1, Quan Wang2, Zhongnan Wang1, Yongdong Zhang1, Zhendong Mao1, 1University of Science and Technology of China, 2Beijing University of Posts and Telecommunication
1 Introduction
@@
Large language models (LLMs) are renowned for their text understanding and generation capabilities (Brown et al., 2020; Achiam et al., 2023; Touvron et al., 2023; Radford et al., 2019). Despite their widespread use, LLMs often produce factual errors, including hallucinations and outdated information (Ji et al., 2023; Wang et al., 2023; Tam et al., 2023). Retraining or fine-tuning to update the model is expensive and time-consuming. Therefore, model editing techniques, which modify specific knowledge within LLMs with low resources, are gaining increasing attention (Yao et al., 2023; Mitchell et al., 2022; Meng et al., 2022). In practice, evolving world knowledge necessitates repeated model edits over time, which is known as lifelong model editing (Hartvigsen et al., 2024; Yu et al., 2024).
@@
The most intuitive way to implement lifelong editing is to perform model editing methods successively for multiple times. However, most model editing methods are designed for one-time use (Meng et al., 2022; Mitchell et al., 2022). Repeated using them causes LLMs to forget previous edits and pre-training data, significantly reducing their edit reliability and general ability on downstream tasks (Yao et al., 2023; Gu et al., 2024; Yang et al., 2024; Gupta et al., 2024; Lin et al., 2024; Huang et al., 2022). Recently, some methods have been tailored for lifelong model editing (Hartvigsen et al., 2024; Yu et al., 2024). These methods freeze the original LLM parameters and incorporate additional adapters to modify the model. They cluster the input data and assign a specific adapter to each cluster, maintaining discrete data-adapter mappings. This approach enables the model to manage different knowledge with independent parameters, preventing interference between different edits. As a result, it ensures high reliability after sequential edits and outperforms other techniques.
@@
However, these key-value mapping methods exhibit poor robustness and struggle with semantically equivalent inputs (Tian et al., 2024; Lin et al., 2024). Due to the inherent discreteness of their data-adapter mappings, data points on opposite sides of a cluster boundary will map to entirely different adapters. Unfortunately, their clustering relies on manually set distance metrics and hyperparameters, resulting in inaccurate cluster boundaries. Semantically equivalent data with slight variations (e.g., rephrased sentences) could fall outside the appropriate cluster and are assigned incorrect adapters. Consequently, these methods are not robust and prone to errors with rephrased edit data.
@@
To address the robustness issues in previous discrete mapping methods, we propose an adaptive and continuous learning method named ELDER (Enhancing Lifelong moDel Editing with mixtuRe of Low-Rank Adapter). ELDER facilitates precise and successive updates to model knowledge by utilizing a router network to integrate multiple Low-Rank Adapters (LoRAs) Hu et al. (2021), akin to the Mixture-of-Experts (MoE) structure (Jacobs et al., 1991; Jordan and Jacobs, 1994). Unlike previous methods, it adaptively generates LoRA weights through end-to-end learning instead of manually set distance metrics, and produces a weighted combination of top-$k$ LoRA outputs. For different edits, ELDER dynamically adjusts LoRA weights to produce varied adapter allocations based on edit semantics, ensuring that semantically equivalent inputs are assigned similar allocations to generate consistent model responses. This approach maintains a continuous and smooth data-adapter connection, eliminating the risk of using completely different adapter parameters due to slight input variations posed by discrete cluster boundaries. Thus, it ensures robust performance. To the best of our knowledge, this is the first work to employ a mixture-of-LoRA structure for model editing. Another advantage of ELDER lies in its scalability. Unlike discrete mapping methods, which require new parameters for each modification, ELDER manages knowledge modifications seamlessly by various LoRA combinations rather than independent adapters. Therefore, this approach avoids the need for additional parameters with each successive edit, allowing for scalability to longer editing sequences.
@@
To further enhance ELDER, we propose a guided loss function to align adapter allocation with data semantics and a deferral mechanism to retain general capabilities in post-edit models. Firstly, in the training data, semantically equivalent edits are preset with the same LoRA allocation. We design an auxiliary training loss to guide the model in learning these allocations and establishing the association between LoRA allocation and data semantics, thus promoting the model to assign similar LoRA allocations to similar inputs during inference. Moreover, our deferral mechanism identifies whether an input requires editing based on its LoRA allocation. In this way, it concentrates on edit-related features while ignoring irrelevant details like input format to ensure accurate discrimination. For test inputs that differ significantly from edited samples, this mechanism deactivates the mixture-of-LoRA, retaining the model’s original performance on downstream tasks while leveraging LoRAs for specific edits.
@@
Through extensive experiments, we have demonstrated the effectiveness of ELDER. Our experiments are conducted on two popular LLMs, i.e., GPT2-XL(Radford et al., 2019) and LLaMA2-7B (Touvron et al., 2023), using two widely used model editing datasets. Results indicate that ELDER achieves better editing performance and enhances the robustness of rephrased edits by improving the editing generalization by over 10% higher than existing methods. It is also superior in reliably maintaining previous edits. Furthermore, we show that ELDER retains most of post-editing LLM’s abilities on downstream general tasks, significantly surpassing most existing methods.
2 Method
Our proposed ELDER incorporates multiple mixture-of-LoRA modules, which adaptively allocate LoRAs to successive edits and establish a continuous and smooth connection between data and adapters. Moreover, an auxiliary training loss is specially designed for lifelong model editing to assist the model in aligning LoRA allocations with input semantics. During inference, these LoRA allocations help to identify task inputs that do not require editing via a simple yet effective deferral mechanism. Such inputs are processed using the original LLM, thereby preserving the model’s performance on general tasks.
@@ 2 Method
Figure 1: An illustration of processing two different edits with ELDER. The mixture-of-LoRA module is applied to the FC layer at the FFN of the Transformer block. Each edit is routed to top-$k$ LoRAs with the highest scores based on its query vector. This figure takes $k=1$ as an example. The final results are summations of LoRA outputs and outputs of the original FC. Dotted lines denote multiplying LoRA outputs with corresponding weights.
2.1 Problem Formulation
@@
The lifelong model editing task, as described by Hartvigsen et al. (2024), involves continuously editing an initial base LLM, $f_{base}$, without diminishing its overall performance or negating the corrections made by prior edits. Let $D_{edit}=\{e_{1},\dots,e_{n}\}$ denote $n$ successive edits applied to the model, where each edit $e_{i}=(x_{i},y_{i})$. The primary objective is to obtain a post-edit model $f_{n}$ that correctly maps an input $x$ to its corresponding prediction $y$, i.e., $f_{n}(x_{i})=y_{i}$. Furthermore, $f_{n}$ should be capable of predicting the equivalent neighbor $N(e_{i})$ (Yao et al., 2023), such as rephrased sentences, ensuring $f_{n}(x_{i}^{\prime})=y_{i}^{\prime}$ for all $(x_{i}^{\prime},y_{i}^{\prime})\in N(e_{i})$. Additionally, $f_{n}$ must retain the performance on its original test data for practical usability(Hartvigsen et al., 2024). Given $k$ representative tasks $t_{1},\dots,t_{k}$ and their respective metrics $m_{1},\dots,m_{k}$, $f_{n}$ should preserve $f_{base}$’s capabilities across these tasks, i.e., $m_{j}(t_{j},f_{n})=m_{j}(t_{j},f_{base}),\forall j\in\{1,\dots,k\}$.
2.2 Mixture-of-LoRA Structure
@@
ELDER employs mixture-of-LoRA modules to enhance robustness and establish a smooth association between sequential data and adapters. Each mixture-of-LoRA module comprises $N$ LoRAs and a router network, as shown in Figure 1. The router network, implemented as a fully connected (FC) layer, takes a text sequence’s query vector as input and routes the entire sequence to the top-$k$ LoRAs selected from a pool of $N$ LoRAs. All tokens in the same sequence are routed to the same LoRA allocation, ensuring equal treatment of the entire knowledge. Following prior works (Hartvigsen et al., 2024; Yu et al., 2024), we define the query vector as the hidden representation of the last token in the input sequence, denoted as $\textbf{x}\in\mathbb{R}^{d}$. The router network generates scores $\mathbf{s}(\mathbf{x})$ for each LoRA by normalizing the projection of x using a softmax distribution over the available $N$ LoRAs at that layer, which are given by:
@@
We further apply top-$k$ gating (Shazeer et al., 2016), selecting $k$ LoRAs with the highest scores for routing. Each selected LoRA module generates a product matrix $\Delta\mathbf{W}_{i}=\mathbf{B}_{i}\mathbf{A}_{i}$, where $\mathbf{B}_{i}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}_{i}\in\mathbb{R}^{r\times k}$ are two low-rank matrices in the $i$-th LoRA module. The updated matrix $\Delta\mathbf{W}$ is obtained by computing the linearly weighted combination of the selected LoRA matrices based on the score values,
@@
Here, $\mathbf{W}_{0}$ is the FC layer’s original weight matrix, which remains frozen during training. The modified FC computation is applied to all tokens in the given input sequence. Furthermore, as illustrated in the figure, ELDER can handle more edits through adapter combinations without requiring additional learnable parameters for each new edit. In contrast, previous discrete mapping methods increase the parameter count linearly with each new edit, implying scalability issues.
2.3 Guided Loss
We propose an auxiliary training loss to assist the model in assigning LoRA allocations based on input semantics. Specifically, we first pre-assign a LoRA allocation via random generation for each edit in the training data, ensuring that identically labeled edits receive the same allocation distinct from others. Subsequently, our proposed training loss guides the router network in learning to assign these allocations. This approach promotes the model to associate input semantics with various adapter allocations, thereby cultivating the adaptability to assign similar allocations to semantically equivalent inputs during inference.
Formally, we guide the model with a new loss, $L_{guide}$, by maximizing the probability of selecting the pre-assigned LoRAs. Suppose the model has $L$ layers with mixture-of-LoRA modules. Each produced allocation $\mathcal{A}$ should contain $L\times k$ indices, indicating selected adapters from a total $L\times N$ available LoRAs. Therefore, the loss function takes the form of:
@@ 2.3 Guided Loss
Although previous MoE methods (Fedus et al., 2021; Dai et al., 2022b) improve the token-routing scheme by balancing the usage load of parallel modules to achieve uniform distribution, they present a challenge in lifelong model editing. In this context, sequential edits occur sparsely, and all tokens in a single edit should apply the same routing scheme rather than being routed separately, as they represent the same knowledge. Thus, each training batch has fewer samples, making it difficult to calculate the average usage load accurately. This disparity means that batch measurements do not accurately reflect real usage loads, rendering the optimization for a uniform distribution inapplicable to our task.
2.4 Deferral Mechanism
Algorithm 1 Inference with deferral.
@@ 2.4 Deferral Mechanism
3:Input: Edit allocation codes $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$
@@
During inference, we aim to preserve the original capabilities of LLMs on general tasks to ensure their practical application. To this end, we implement a simple yet effective deferral mechanism to identify inputs that do not involve edited knowledge, referred to as task inputs. The proposed mechanism is based on the LoRA allocation since it contains edit-related features within the input, excluding the impact of irrelevant details like input format as described in Section 2.3. After distinguishing these non-edited inputs, we deactivate the mixture-of-LoRA module and process them directly with the original model. This mechanism ensures that task inputs yield the same output as they would from the initial model.
@@
where $(i,j)$ denotes the indices of the $j$-th LoRA at the $i$-th mixture-of-LoRA layer. $\mathbf{c}$ is computed using all router networks before the first mixture-of-LoRA (line 6) to discriminate model inputs in advance for efficient editing. The allocation codes $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$ for all edited samples are precomputed and stored during training. The nearest distance between $\mathbf{c}$ and all items in $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$ is then calculated using the Hamming distance (line 7). Inputs whose nearest distance exceeds a threshold $\epsilon$ are identified as task inputs and are processed with the preserved original parameters (line 9).
3 Experiments
3.1 Experimental Setup
3.1.1 Baselines
We compare our proposed method, ELDER, against several advanced baseline methods. We first select two methods tailored for the lifelong editing setup. They preserve the original model and modify it with discrete data-adapter mappings. GRACE (Hartvigsen et al., 2024) uses representation vectors as adapters, writing them into the pre-trained model’s latent space based on specific data samples. MELO (Yu et al., 2024) builds on the same foundation framework but uses LoRAs as adapters. Additionally, we compare with Defer, a model editor based on SERAC (Mitchell et al., 2022) and adapted for lifelong editing by Hartvigsen et al. (2024). ROME (Meng et al., 2022), a state-of-the-art one-time model editing method, is also selected as a baseline. It locates the edit area in GPTs via casual tracing and updates relevant weights. Beyond these model editing techniques, we compare with FT-L (Meng et al., 2022), which fine-tunes the layers identified by ROME.
3.1.2 Metrics
We apply several evaluation metrics consistent with those described in previous works (Yao et al., 2023; Hartvigsen et al., 2024; Yu et al., 2024).
@@ 3.1.2 Metrics
Following previous works in lifelong model editing (Hartvigsen et al., 2024; Yu et al., 2024), we check how well $f_{n}$ retains its original performance during inference. We evaluate $f_{n}$ on $k$ representative general tasks and take the average result number as its test retention on these tasks:
@@
3.1.3 Datasets
We evaluate the reliability and generalization of ELDER on lifelong model editing with two widely used model editing datasets for training and evaluation. The first one, ZsRE (Levy et al., 2017), is a zero-shot relation extraction dataset that uses question rephrasing generated through back-translation as the rephrased edit data. The second dataset, CounterFact (Meng et al., 2022), is a more challenging dataset. It comprises counterfactual statements initially receiving low factuality scores. We adopt both datasets to the lifelong model editing setting by extracting a sequence of 1000 editing samples with their rephrasings for our main experiments, following the methodologies outlined in (Hartvigsen et al., 2024) and (Yu et al., 2024). Further details are in Appendix B.1.

### v2

# ELDER: Enhancing Lifelong Model Editing with Mixture-of-LoRA

@@ Abstract
Large language models (LLMs) require model editing to efficiently update specific knowledge within them and avoid factual errors. Most model editing methods are solely designed for single-time use and result in a significant forgetting effect in lifelong editing scenarios, where sequential edits are conducted over time. Previous approaches manage sequential edits by freezing original parameters and discretely allocating new parameters for each knowledge update. However, these methods lack robustness to minor input variations due to the discrete mapping between data and parameters. To overcome this challenge, we propose ELDER, a novel approach to create a continuous association between data and adapters. ELDER integrates multiple LoRAs through a router network and is trained to establish a smooth data-adapter association, thereby enhancing the edit robustness and generalization of semantically equivalent inputs. To ensure inputs containing the same knowledge will be processed by the same LoRAs, we design a novel loss to guide the model link LoRA allocations with edit knowledge. Furthermore, we propose a deferral mechanism to retain the original LLM capabilities post-edit. Extensive experiments on GPT-2 XL and LLaMA2-7B demonstrate that ELDER effectively edits models in the lifelong setting, outperforming eight baselines while exhibiting strong scalability and preserving LLMs’ general abilities on downstream tasks.
Introduction
@@
Large language models (LLMs) are renowned for their text understanding and generation capabilities (Brown et al. 2020; Achiam et al. 2023; Touvron et al. 2023; Radford et al. 2019). Despite their widespread use, LLMs often produce factual errors, including hallucinations and outdated information (Ji et al. 2023; Wang et al. 2023; Tam et al. 2023). Retraining or fine-tuning to update the model is expensive and time-consuming. Therefore, model editing techniques, which modify specific knowledge within LLMs with low resources, are gaining increasing attention (Yao et al. 2023; Mitchell et al. 2022; Meng et al. 2022). In practice, evolving world knowledge necessitates repeated model edits over time, which is known as lifelong model editing (Hartvigsen et al. 2024; Yu et al. 2024).
@@
The most intuitive way to implement lifelong editing is to perform model editing methods successively for multiple times. However, most model editing methods are designed for one-time use (Meng et al. 2022; Mitchell et al. 2022). Repeated using them causes LLMs to forget previous edits and pre-training data, significantly reducing their edit reliability and general ability on downstream tasks (Yao et al. 2023; Gu et al. 2024; Yang et al. 2024b; Gupta, Rao, and Anumanchipalli 2024; Lin et al. 2024; Huang et al. 2022). Recently, a class of representative methods has been developed specifically for lifelong model editing (Hartvigsen et al. 2024; Yu et al. 2024). These methods freeze the original LLM parameters and incorporate additional adapters to modify the model. They cluster the input data and assign a specific adapter to each cluster, maintaining discrete data-adapter mappings. This approach enables the model to manage different knowledge with independent parameters, preventing interference between different edits. As a result, it ensures high reliability after sequential edits and outperforms other techniques.
@@
However, these key-value mapping methods exhibit poor robustness and struggle with semantically equivalent inputs (Tian et al. 2024; Lin et al. 2024). Due to the inherent discreteness of their data-adapter mappings, data points on opposite sides of a cluster boundary will map to entirely different adapters. Unfortunately, their clustering relies on manually set distance metrics and hyperparameters, resulting in inaccurate cluster boundaries. Semantically equivalent data with slight variations (e.g., rephrased sentences) could fall outside the appropriate cluster and are assigned incorrect adapters. Consequently, these methods are not robust and prone to errors with rephrased edit data.
@@
To address the robustness issues in previous discrete mapping methods, we propose a novel approach to create an adaptive and continuous association between factual knowledge and adapter parameters, named ELDER (Enhancing Lifelong moDel Editing with mixtuRe-of-LoRA). ELDER successively updates model knowledge by utilizing a router network to integrate multiple LoRAs (Hu et al. 2021), akin to the Mixture-of-Experts (MoE) structure (Jacobs et al. 1991; Jordan and Jacobs 1994). Unlike previous lifelong model editing methods, it adaptively generates LoRA weights through end-to-end learning instead of manually-set distance metrics, and produces a weighted combination of top-$k$ LoRA outputs. For different edits, ELDER dynamically adjusts LoRA weights to produce varied adapter allocations based on edit semantics, ensuring that semantically equivalent inputs are assigned similar allocations to generate consistent model responses. Thus, it ensures robust performance. Another advantage of ELDER lies in its scalability. Unlike discrete mapping methods, which require new parameters for each modification, ELDER manages knowledge modifications seamlessly by various LoRA combinations rather than independent adapters. Therefore, this approach avoids the need for additional parameters with each successive edit, allowing for scalability to longer editing sequences. To the best of our knowledge, this is the first work to employ a mixture-of-LoRA structure for model editing.
@@
Beyond employing Mixture-of-LoRA, we also introduce two novel techniques designed to address specific challenges in the lifelong editing task. Specifically, we propose a guided loss function to align adapter allocation with edit knowledge and a deferral mechanism to retain general capabilities in post-edit models. Firstly, in the training data, semantically equivalent edits are preset with the same LoRA allocation. We design a novel training loss to guide the model in learning these allocations and establishing the association between LoRA allocation and data semantics, thus promoting the model to assign similar LoRA allocations to similar inputs during inference. Moreover, our deferral mechanism identifies whether an input requires editing based on its LoRA allocation. In this way, it concentrates on edit-related features while ignoring irrelevant details like input format to ensure accurate discrimination. For test inputs that differ significantly from edited samples, this mechanism deactivates the mixture-of-LoRA, retaining the model’s original performance on downstream tasks while leveraging LoRAs for specific edits.
@@
Through extensive experiments, we have demonstrated the effectiveness of ELDER. Our experiments are conducted on two popular LLMs, i.e., GPT2-XL(Radford et al. 2019) and LLaMA2-7B (Touvron et al. 2023), with two widely used model editing datasets, ZsRE (Levy et al. 2017) and CounterFact (Meng et al. 2022). Results indicate that ELDER achieves better editing performance and enhances the robustness of rephrased edits by improving the editing generalization by over 10% higher than eight baselines. It is also superior in reliably maintaining previous edits. Furthermore, we show that ELDER retains most of post-editing LLM’s abilities on downstream general tasks, significantly surpassing most existing methods.
Related Works
Model Editing
Model editing aims to accurately and efficiently modify knowledge in deep neural networks(Sinitsin et al. 2019), especially language models(Yao et al. 2023). Meta-learning-based methods MEND(Mitchell et al. 2021) and KE(De Cao, Aziz, and Titov 2021) train an extra hypernetwork to learn changes in the base model. KN(Dai et al. 2022a) attributes the neuron that embodies the knowledge and updates these neurons. ROME(Meng et al. 2022) locates the edit area by casual analysis and modifies the entire weight matrix. SERAC(Mitchell et al. 2022) trains a scope classifier to identify inputs that need to be edited, then uses a counterfactual model to process these inputs separately from the original model.
However, the above methods only consider static edits, i.e., modifying the model a single time. Huang et al. (2022) proposed a sequential editing approach by adding one neuron for each edit sample. Nevertheless, this method relies on large sets of unrelated inputs and is slow due to the need to train neurons for each edit (Yao et al. 2023). Most similar works to ours are GRACE (Hartvigsen et al. 2024) and MELO(Yu et al. 2024), which discretely maps different adapters to successive edits.
Mixture-of-Experts (MoE)
Jacobs et al. (1991); Jordan and Jacobs (1994) introduced MoE models to compute different examples with independent expert modules. Shazeer et al. (2016) extended this concept to large-scale language models with LSTMs. Advances such as GShard (Fedus, Zoph, and Shazeer 2021), Switch Transformer (Fedus, Zoph, and Shazeer 2021), and BASE Layer Lewis, Yao, and Ruder (2021) have refined routing input tokens to experts. Hash Layer (Roller et al. 2021) uses a pre-defined token-level hash table for token-to-expert assignment. StableMoE (Dai et al. 2022b) addresses routing fluctuation by training the router network first and then freezing it for further model training. In contrast, our approach guides the router network to learn a pre-set sample-to-adapter assignment, effectively handling unseen but equivalent inputs.
Mixture-of-LoRAs
LoRA (Hu et al. 2021) uses low-rank matrices to update the LLMs, and has become a popular parameter-efficient fine-tuning method. Recently, researchers have been using combinations of multiple LoRAs for further benefits. Huang et al. (2023) proposed composing existing LoRA modules for generalization on new tasks. Zadouri et al. (2023) combined LoRA with token-level mixture-of-experts (Jacobs et al. 1991) and designed a new PEFT module. Dou et al. (2023) split LoRA experts into two separate groups to maintain world knowledge of LLMs during instruction tuning. MoRAL (Yang et al. 2024a) augments LoRA with MoE to achieve both multi-task and fine-tuning abilities.
Although sharing a similar base structure with previous mixture-of-LoRA research, our approach distinguishes itself from these works in following aspects. First, we focus on the task of editing factual knowledge within the model and demonstrating strong effectiveness, which previous works have not explored. Second, our routing scheme is tailored for model editing. We route the whole sequence altogether to ensure equal treatment of knowledge, unlike previous works that apply different experts to individual tokens. Moreover, we introduce novel techniques to enhance the adaptability of mixture-of-LoRA to the lifelong editing task, including the guided loss function and deferral mechanism.
@@
Figure 1: An illustration of processing two different edits with the mixture-of-LoRA module in ELDER. A mixture-of-LoRA module is applied to the FC layer at the FFN of the Transformer block. Each edit is routed to top-$k$ LoRAs with the highest scores based on its query vector. This figure takes $k=1$ as an example. The final results are summations of LoRA outputs and outputs of the original FC. Dotted lines denote multiplying LoRA outputs with corresponding weights. Training loss and deferral mechanism are omitted in this figure for simplicity.
Proposed Methods
Our proposed ELDER establishes a continuous and smooth connection between data and adapters. It contains multiple mixture-of-LoRA modules, which adaptively allocate LoRAs to each successive edit. Moreover, an auxiliary training loss is specially designed for lifelong model editing to assist the model in aligning LoRA allocations with the input knowledge. During inference, these LoRA allocations help to identify task inputs that do not require editing via a novel deferral mechanism tailored for the lifelong editing task. Such inputs are processed using the original LLM, thereby preserving the model’s performance on general tasks.
Problem Formulation
@@
The lifelong model editing task, as described by Hartvigsen et al. (2024), involves continuously editing an initial base LLM, $f_{base}$, without harming its overall abilities or negating the corrections made by prior edits. Let $D_{edit}=\{e_{1},\dots,e_{n}\}$ denote $n$ successive edits applied to the model, where each edit $e_{i}=(x_{i},y_{i})$. The primary objective is to obtain a post-edit model $f_{n}$ that correctly maps an input $x$ to its corresponding prediction $y$, i.e., $f_{n}(x_{i})=y_{i}$. Furthermore, $f_{n}$ should be capable of predicting the equivalent neighbor $N(e_{i})$ (Yao et al. 2023), such as rephrased sentences, ensuring $f_{n}(x_{i}^{\prime})=y_{i}^{\prime}$ for all $(x_{i}^{\prime},y_{i}^{\prime})\in N(e_{i})$. Additionally, $f_{n}$ must retain the performance on its original test data for practical usability(Hartvigsen et al. 2024). Given $k$ example tasks $t_{1},\dots,t_{k}$ and their respective metrics $m_{1},\dots,m_{k}$, $f_{n}$ should preserve $f_{base}$’s abilities across them, i.e., $m_{j}(t_{j},f_{n})=m_{j}(t_{j},f_{base}),\forall j\in\{1,\dots,k\}$.
Mixture-of-LoRA Structure
@@
ELDER employs mixture-of-LoRA modules as the base structure. Each mixture-of-LoRA module comprises $N$ LoRAs and a router network, as shown in Figure 1. The router network, implemented as a fully connected (FC) layer, takes a text sequence’s query vector as input and routes the entire sequence to the top-$k$ LoRAs selected from a pool of $N$ LoRAs. Notably, we route all tokens in the same instance to one LoRA allocation based on the query vector, ensuring equal treatment of the entire knowledge. Following prior works (Hartvigsen et al. 2024; Yu et al. 2024), we define the query vector as the hidden representation of the last token in the input sequence, denoted as $\textbf{x}\in\mathbb{R}^{d}$. The router network generates scores $\mathbf{s}(\mathbf{x})$ for each LoRA by normalizing the projection of x using a softmax distribution over the available $N$ LoRAs at that layer, which are given by:
@@
We further apply top-$k$ gating (Shazeer et al. 2016), selecting $k$ LoRAs with the highest scores for routing. Each selected LoRA module generates a product matrix $\Delta\mathbf{W}_{i}=\mathbf{B}_{i}\mathbf{A}_{i}$, where $\mathbf{B}_{i}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}_{i}\in\mathbb{R}^{r\times k}$ are two low-rank matrices in the $i$-th LoRA module. The updated matrix $\Delta\mathbf{W}$ is obtained by computing the linearly weighted combination of the selected LoRA matrices based on the score values,
@@
Here, $\mathbf{W}_{0}$ is the FC layer’s original weight matrix, which remains frozen during training. The modified FC computation is applied to all tokens in the given input sequence. Furthermore, as illustrated in the figure, ELDER can handle more edits through adapter combinations without requiring additional learnable parameters for each new edit. In contrast, previous discrete mapping methods increase the parameter count linearly for each new edit, implying scalability issues.
Guided Loss
We propose a novel training loss, guiding the model to allocate LoRAs based on knowledge embedded in each input. This training objective promotes the model to associate input semantics with various adapter allocations, thereby cultivating the adaptability to assign similar allocations to semantically equivalent inputs during inference. Specifically, we first pre-assign a LoRA allocation for each edit in the training data via random generation, ensuring that identically labeled edits receive the same allocation distinct from others. Subsequently, our proposed training loss guides the router network in learning to assign these allocations.
Formally, we guide the model with a new loss, $L_{guide}$, by maximizing the probability of selecting the pre-assigned LoRAs. Suppose the model employs mixture-of-LoRA to $L$ layers. Each produced allocation $\mathcal{A}$ should contain $L\times k$ indices, indicating selected adapters from a total $L\times N$ available LoRAs. Therefore, the loss function takes the form of:
@@
Although previous MoE methods (Fedus, Zoph, and Shazeer 2021; Dai et al. 2022b) improve the token-routing scheme by balancing the usage load of parallel modules to achieve uniform distribution, they present a challenge in lifelong model editing. In this context, sequential edits occur sparsely, and all tokens in a single edit apply the same routing scheme rather than being routed separately, as they represent the same knowledge. Thus, each training batch has fewer samples, making it difficult to calculate the average usage load accurately. This disparity means that batch measurements do not accurately reflect real usage loads, rendering the optimization for a uniform distribution inapplicable to our task.
Deferral Mechanism
Algorithm 1 Inference with deferral mechanism.
@@
3:Input: Allocation codes of all edits $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$
@@
During inference, we aim to preserve the original capabilities of LLMs on general tasks to ensure their practical application. To this end, we design a novel deferral mechanism to identify inputs that do not involve edited knowledge, referred to as task inputs. The proposed mechanism is based on the LoRA allocations, since they contain edit-related features within the input, excluding the impact of irrelevant details like input format as described in Section Guided Loss. After distinguishing these non-edited inputs, we deactivate the mixture-of-LoRA module and process them directly with the original model. This mechanism ensures that task inputs yield the same output as they would from the initial model.
@@
where $(i,j)$ denotes the indices of the $j$-th LoRA at the $i$-th mixture-of-LoRA layer. $\mathbf{c}$ is represented by boolean values, and computed using all router networks before the first mixture-of-LoRA (line 6) to discriminate model inputs in advance for efficient editing. The allocation codes $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$ for all edits are precomputed and stored during the editing stage The nearest distance between $\mathbf{c}$ and all items in $\{\mathbf{c}_{e}^{i}\}_{i=1}^{n}$ is then calculated using the Hamming distance (line 7). This distance computation is highly efficient because all allocation codes are boolean, allowing the operations between them to be performed using simple bitwise operations. Inputs whose nearest distance exceeds a threshold $\epsilon$ are identified as task inputs and are processed with the preserved original parameters (line 9).
Experiments
Experimental Setup
Baselines
We compare our proposed method, ELDER, with eight baselines. We first select two methods tailored for the lifelong editing setup. They preserve the original model and modify it with discrete data-adapter mappings. GRACE (Hartvigsen et al. 2024) uses representation vectors as adapters, writing them into the pre-trained model’s latent space based on specific data samples. MELO (Yu et al. 2024) builds on the same foundation framework but uses LoRAs as adapters. Other baselines include SERAC (Mitchell et al. 2022), which additionally trains a scope classifier and a counterfactual model to alter the model behavior, and T-Patcher (Huang et al. 2022), which adjusts one neuron for each new edit. We also include ROME (Meng et al. 2022), a state-of-the-art one-time model editing method. It locates the edit area in GPTs via casual tracing and updates relevant weights. We further include WilKE(Hu et al. 2024), which modifies different layers for different edits. Beyond these model editing techniques, we compare ELDER with FT-L (Meng et al. 2022), which fine-tunes the layers identified by ROME, and with the standard LoRA (Hu et al. 2021).
Metrics
We apply three metrics to evaluate both the editing performance and the preservation of post-edit model capability. These metrics align with those reported in previous works (Meng et al. 2022; Gu et al. 2024; Yao et al. 2023).
@@
We aim to evaluate the side effects of model editing on the capabilities of LLMs. Although a previous metric, Locality, is designed for this purpose, it is limited in scope, focusing on a narrow data range and simple QA task, which is insufficient to assess the full functionality of LLMs, which can lead to inflated results and fails to assess the full functionality of LLMs (Yang et al. 2024b; Gupta, Rao, and Anumanchipalli 2024). To achieve a more comprehensive evaluation, we extend the assessment to a broader range of tasks, following the approach of Gu et al. (2024). Specifically, we measure how well $f_{n}$ retains its original performance during inference across $k$ representative general tasks. The average result is then used to determine its retention on these tasks:
@@
Datasets
We evaluate the reliability and generalization of ELDER on lifelong model editing with two widely used model editing datasets for training and evaluation. The first one, ZsRE (Levy et al. 2017), is a zero-shot relation extraction dataset that uses question rephrasing generated through back-translation as the rephrased edit data. The second dataset, CounterFact (Meng et al. 2022), is a more challenging dataset. It comprises counterfactual statements initially receiving low factuality scores. We adopt both datasets to the lifelong model editing setting by extracting a sequence of 1000 editing samples with their rephrasings for our main experiments, following the methodologies outlined in (Hartvigsen et al. 2024) and (Yu et al. 2024). Further details are in the technical appendix.